# Multi-Relational Algebra and Its Applications to Data Insights

### v2

@@ Abstract
A range of data insight analytical tasks involves analyzing a large set of tables of different schemas, possibly induced by various groupings, to find salient patterns. In particular, such analyses are about many-to-many transformations of tables, while the classic relational algebra is about one-to-one or many-to-one transformations. This paper presents Multi-Relational Algebra, which extends relational algebra for such transformations and their compositions. Multi-Relational Algebra introduces MultiRelation to model of a set of tables of different schemas. Importantly, while the information unit in Relational Algebra is a tuple, the information unit in Multi-Relational Algebra is a slice, which formally is a pair $(r,X)$ where $r$ is a (region) tuple, and $X$ is a (feature) table [^1] In this paper, “table” and “relation” are interchangeable.. Multi-Relational Algebra introduces three new fundamental algebraic operators, MultiSelect, MultiProject, and MultiJoin, which lift their counterparts Select, Project, and Join to transform MultiRelation to MultiRelation. Through various examples, we show that multi-relational algebra can effortlessly express many complex analytic problems, some of which are traditionally considered out of scope for relational analytics. We have implemented and deployed a service for multi-relational analytics. Due to a unified logical design, we are able to conduct systematic optimization for a variety of seemingly different tasks. Our service has garnered interest from over a hundred internal teams who have developed data-insight applications using it, and serves millions of operators daily.
@@ 1 Introduction
Multi-Relational Algebra. We describe (Section 2) Multi-Relational Algebra, which extends the Relational Algebra. Specifically, we define MultiRelation which models a set of slices, and define three new algebraic operators, MultiSelect, MultiProject, and MultiJoin, which lift their classic counterparts, Select, Project, Join, to transform MultiRelation to MultiRelation. Multi-Relational Algebra significantly broadens the analytical scope of relational analytics.
Applications of Multi-Relational Algebra. We have built a service for executing a pipe of multi-relational operators. Our service has garnered interest from over a hundred internal teams who have developed data-insight applications using it, and serves millions of operators daily (Figure 9). Section 4 reports some major applications.
@@ 2.1 Data Model
A multi-relational block, or simply block, is a pair of schema $(\Gamma,\operatorname{{\mathcal{}F}})$ where $\Gamma\cap\operatorname{{\mathcal{}F}}=\emptyset$.
@@ 3 Multi-Relational Analytics
It should not be a surprise that multi-relational operators with proper slice transformations can solve a wide-variety of problems. For example, MultiSelect with some trivial slice transformations capture the following two previous work:
@@
  • AutoSlicer (Liu et al. 2022). [LRC22] proposed AutoSlicer to help understand deeper model performance in complex data. Their “search space” is exactly the region space in CRAWL and the analysis on slices can be easily captured by a region analysis model.
@@ 3.1 Frequent Itemset Mining
Dividing these two thus computes the support for a pattern. Several remarks: (1) Note that FIM is not traditionally considered a relational analytics problem, yet this example shows that it can be effortlessly captured by Multi-Relational Algebra, a natural extension of Relational Algebra. In particular, this example shows the power of having a slice transformation analyzing multiple slices (in this case, a region slice and a reference slice). (2) There are many optimizations we can do to evaluate this MultiSelect efficiently for this particular slice transformation. For example, the classic Apriori algorithm [Apr] can be implemented as a breadth-first (over regions) evaluation of the MultiSelect. A detailed discussion of this is however beyond the scope of this paper, which we leave as future work.
@@ 3.5 Recurrent CRAWL
4 Applications
We have built a service for executing a pipe of multi-relational algebra operators. Our service has attracted more than a hundred internal teams to build data insights applications with it. Figure 9 plots the daily usage of the service since Sept, 2023.
@@ 4 Applications
The multi-relational algebra pipes service provides high-level abstractions that benefit all these applications:
Our approach drastically simplifies system construction. For instance, in experimentation analysis platforms, teams often develop custom slicing-analysis kernels, each with tens of thousands of lines of code, typically built on Beam for its flexibility over SQL. Our multi-relational algebra pipes streamline these kernels into a concise pipeline of operators, enhancing interoperability with SQL due to their shared relational data foundation. This results in significant simplification of both experimentation analysis and integration with relational databases.
Enabling Advanced Capabilities. The composability of multi-relational operators has unlocked new and powerful functionalities. For example, in system monitoring, techniques like “recurrent crawl,” 3.5 built on these operators, have achieved superior precision and recall in anomaly detection compared to traditional statistical methods. This demonstrates the ability of multi-relational operators to express complex analysis patterns that would be difficult to achieve otherwise.
Enhanced Capabilities, Reduced Costs. The flexibility of slice transformations enables seamless integration of the Python Data Science stack with multi-relational processing. This allows for advanced analysis, such as quickly identifying revenue anomalies and their correlation with campaign budget allocation, leading to significant revenue gains in some cases. Additionally, migrating ad-hoc experimentation analysis systems to multi-relational pipes not only expands their capabilities to analyze finer-grained regions but also surprisingly results in improved latency, scalability, and reduced resource costs. This efficiency stems from the unified logical design of multi-relational algebra, enabling systematic optimizations across diverse analytical tasks.
@@ Appendix A Emulating WINDOW functions
Appendix B Extended operators
@@
B.1 MultiGroupBy
In the RA, GroupBy partitions a relation into groups, and then uses an aggregation function to condense each group into a tuple, and finally we coalesce these tuples back into one relation. For BiRA, the situation is similar, we partition a bilevel relation into groups, where each group is a bilevel relation. Then we transform each group into another bilevel relation. Finally, we coalesce these bilevel relations into one by flattening the group key into the regions of each bilevel relation, and take union of all region-feature pairs.
@@ B.1 MultiGroupBy
BiGroupBy is an operator as follows:
BiGroupBy(
br,
@@
    We partition br using by_attributes.
@@
    For each group, we apply bilevel_transform to transform each bilevel relation to another.
@@
    We flatten the group keys into the regions of t bilevel relation, and take union of all region-feature pairs.
B.2 CoMultiGroupBy
Consider two bilevel relations br1 and br2, and we groupby for each on Country. Now, we want the bilevel relation in br1 under Country=US to co-occur with the bilevel relation in br2 under Country=US. Such a need of co-occurrence is not new: this also appears in Apache Beam CoGroupByKey [CoG]. However, unlike Beam which only needs to work with a data model of general dictionaries, here we must eventually output one bilevel relation. Suppose we use br1{Country=US} to denote the group Country=US after groupby, then co-groupby will put this together with br2{Country=US}, so we have a pair of bilevel relations (br1{Country=US}, br2{Country=US}). Then, we can use a BiJoin to condense this pair of bilevel relations back into one. After this, we are then in the situation of BiGroupBy, which we know how to handle.
@@ B.2 CoMultiGroupBy
CoBiGroupBy is an operator as follows:
CoBiGroupBy(
br1,
br2,
@@
    We partition br1 and br2 by by_attributes to form groups, where each group is a bilevel relation.
@@
    For each group key K, we have a bilevel relation br1{K}, and another bilevel relation br2{K}.
@@
    We join these two bilevel relations into one, via bijoin(br1{K}, br2{K}).
After the above steps, we are in the same situation as BiGroupBy, which we handle accordingly.

### v3

@@ Abstract
A range of data insight analytical tasks involves analyzing a large set of tables of different schemas, possibly induced by various groupings, to find salient patterns. This paper presents Multi-Relational Algebra, an extension of the classic Relational Algebra, to facilitate such transformations and their compositions. Multi-Relational Algebra has two main characteristics: (1) Information Unit. The information unit is a slice $(r,X)$, where $r$ is a (region) tuple, and $X$ is a (feature) table [^1] In this paper, “table” and “relation” are interchangeable.. Specifically, a slice can encompass multiple columns, which surpasses the information unit of “a single tuple” or “a group of tuples of one column” in the classic relational algebra, (2) Schema Flexibility. Slices can have varying schemas, not constrained to a single schema. This flexibility further expands the expressive power of the algebra. Through various examples, we show that multi-relational algebra can effortlessly express many complex analytic problems, some of which are beyond the scope of traditional relational analytics. We have implemented and deployed a service for multi-relational analytics. Due to a unified logical design, we are able to conduct systematic optimization for a variety of seemingly different tasks. Our service has garnered interest from numerous internal teams who have developed data-insight applications using it, and serves millions of operators daily.
@@ 1 Introduction
Multi-Relational Algebra. Section 2 describes Multi-Relational Algebra, which extends Relational Algebra. Specifically, this algebra introduces MultiRelation for modeling a set of slices, and three new operators, MultiSelect, MultiProject, and MultiJoin, which lift their classic counterparts, Select, Project, Join, to transform MultiRelation to MultiRelation. Multi-Relational Algebra significantly broadens the analytical scope of relational analytics. For example, it captures two prior work, DIFF [ABD+18] and AutoSlicer [LRC22], both previously considered “beyond relational algebra.”
Multi-Relational Analytics. We have built a service for constructing and evaluating multi-relational analytics. Our service has garnered interest from numerous internal teams who have developed applications using it, and serves millions of operators daily (Figure 9). Section 4 gives an overview of the service, and reports main application categories.
@@ 2.1 Data Model
A multi-relational block, or simply block, is a pair of schemas $(\Gamma,\operatorname{{\mathcal{}F}})$, where $\Gamma\cap\operatorname{{\mathcal{}F}}=\emptyset$.
@@ 3 Multi-Relational Analytics
Multi-relational operators with slice transformations capture a wide-variety of problems, some of which are beyond the scope of traditional relational analytics. This power stems from two aforementioned characteristics of this algebra: (1) Information Unit. The information unit is a slice $(r,X)$, which can encompass multiple columns. This surpasses the information unit of “a single tuple” or “a group of tuples of one column” in the classic relational algebra, (2) Schema Flexibility. Slices can have varying schemas, not constrained to a single schema. This flexibility further expands the expressive power of the algebra.
As initial illustrations, MultiSelect combined with simple slice transformations can express the functionality of two prior works, DIFF [ABD+18] and AutoSlicer [LRC22], both previously considered “beyond relational algebra.” DIFF, introduced as a novel relational operator, encompasses the capabilities of various ad-hoc data explanation engines. AutoSlicer, implemented using Apache Beam, serves in machine learning model validation.
@@
  • AutoSlicer (Liu et al. 2022). [LRC22] proposed AutoSlicer to help understand deeper model performance in complex data. Their “search space” is exactly the region space in CRAWL and the analysis on slices can be easily captured by a 1-slice transformation (Definition 6).
@@ 3.1 Frequent Itemset Mining
Dividing these two thus computes the support for a pattern.
Remark 1.
Several remarks:
  1. 1.
    Note that FIM is not traditionally considered a relational analytics problem, yet this example shows that it can be effortlessly captured by Multi-Relational Algebra, a natural extension of Relational Algebra. In particular, this example shows the power of having a slice transformation analyzing multiple slices (in this case, a region slice and a reference slice).
  2. 2.
    There are many optimizations we can do to evaluate this MultiSelect efficiently for this particular slice transformation. For example, the classic Apriori algorithm [Apr] can be implemented as a breadth-first (over regions) evaluation of the MultiSelect. We briefly discuss this optimization in Section 4.2.
  3. 3.
    Astute readers may argue that our approach requires encoding transactions in a very wide table, with most of the table 00 (the number of columns is the number of distinct items), and it is wasteful to do this encoding. There is an elegant way to avoid such waste by exploiting the schema of a multi-relation, which we discuss in Section B.
@@ 3.5 Recurrent CRAWL
4 Multi-Relational Pipe and Applications
We have built a service for executing a pipe of multi-relational algebra operators. Our service has attracted more than a hundred internal teams to build data insights applications with it. Figure 9 plots the daily usage of the service since Sept, 2023, which has a daily peak of serving about 12 million operators. In this section we give a brief overview of our pipe service and its applications. Specifically, Section 4.1 presents an end-to-end example of a multi-relational pipe. Then, Section 4.2 briefly describes main optimizations for evaluating MultiSelect. Finally, we discuss main application categories in Section 4.3.
@@ 4 Multi-Relational Pipe and Applications
4.1 An example multi-relational pipe
Consider a hypothetical scenario where we deploy a service on different devices and browsers, and we collect aggregated Cost and Clicks for each device-browser combinations. Figure 10 gives a hypothetical data. Suppose that an analyst wants to analyze the change of CostPerClick (Cpc) by comparing two set of data points: the data points where is_test=True, and the data point where is_test=False. The aggregated Cpc (SUM(Cost)/SUM(Clicks)) changes significantly: For is_test=True, the overall Cpc is about 15400.15, and for is_test=False, the overall Cpc is 56.25.
Figure 10: pipe_df. A base table for finding insights. For is_test=True, CostPerClick = (100 + 100 + 2 + 200,000)/(1 + 1 + 1 + 10) $\approx$ 15400.15. For is_test=False, CostPerClick = (100 + 100 + 300 + 400)/(4 + 4 + 4 + 4) = 56.25. We want to find regions in Device and Browser that contribute to this change.
The first question. As a first question, the analyst wants to know which regions in Device and Browser are the main contributors. To answer this question, we first filter rows where the metrics are NULL, and then do a CRAWL on the regions in Device and Browser (which contains 3 region spaces, [Device], [Browser], [Device, Browser]) and use density Aumann-Shapley attribution (described in Section 3.2) to analyze each region. This is expressed as a pipe in Listing 1. The ‘density_analysis’ computes the metric aumann_shapley_attribution, which we filter by ‘> .5’. The results of the first pipe are shown in Figure 11. We have found the two rows, row-2 and row-3, as the main contributors. This is encoded in two regions, (Device=Pixel) which corresponds to both row-2 and row-3, and (Device=Pixel, Browser=Edge), which corresponds to row-3.
density_analysis = DensityDifferentiation(
mass_column=’Cost’,
volume_column=’Clicks’,
is_test_column=’is_test’,
results = (
pipe_df
| SqlTransform(’SELECT * FROM pipe WHERE Cost IS NOT NULL’)
| CreateMultiTable(all_region_columns=[’Browser’, ’Device’, ’is_test’],)
| Crawl(
region_spaces=ALL_SPACES([’Browser’, ’Device’]),
slice_transformations=[density_analysis],
predicates=[aumann_shapley_attribution > 0.5,])
Listing 1: A first pipe for finding important regions that contribute to a change. The slice transformation uses the density aumann-shapley attribution described in Section 3.2.
Figure 11: Resulting of executing pipe 1. As expected (from eyeballing the data since it is small enough), we find row-2 and row-3 in table 10 as the main contributors, where row-2 drags the Cpc to be smaller, while row-3 drags the Cpc to be bigger. The region (Device=Pixel,) contains both row-2 and row-3, which have the biggest influence, the second region (Device=Pixel, Browser=Edge) contains only row-3, which is the biggest contributor to make the Cpc bigger for is_test=True.
The second question. Note that Device=Pixel appears in both of the result regions, while Browser=Edge appears in only one. It is a common interest of our users to find frequent attribute patterns. This can be achieved using a second pipe, by crawling on the flattened results of the first analysis, using a frequent itemset analysis (as described in Section 3.1). This is expressed in Listing 2. The results of this pipe are shown in Figure 12: As expected, (Device=Pixel) is found as the most frequent pattern (support=1.0; ‘support’ of a pattern is defined in the first paragraph of Section 3.1).
fim_analysis = FrequentItemsetAnalysis(
count_column=’normalized_aumann_shapley_attribution’,
results2 = (
results
| FlattenMultiRelation()
| CreateMultiRelation(all_region_columns=[’Browser’, ’Device’])
| Crawl(
region_spaces=ALL_SPACES([’Browser’, ’Device’]),
slice_transformations=[fim_analysis],
Listing 2: A second pipe for finding frequent attribute patterns in the results of pipe 1. The slice transformation uses the frequent itemset slice transformation described in Section 3.1.
Figure 12: Resulting of executing pipe 2. As expected, we find the Device=Pixel to be a frequent pattern with support 1, since it appears in all the regions in Figure 11.
4.2 Optimizing MultiSelect with Apriori Property
As we mentioned earlier, MultiSelect is not just brute-force, since evaluation of the slice transformations can be correlated between region spaces. Further, a unified logical modeling of MutltiSelect allows us to optimize for a variety of seemingly different analytical tasks. Some of these optimizations are beyond the scope of classical relational algebra, which is designed for processing set of tuples rather than set of slices. In this section we give a main example: Optimizing MultiSelect with Apriori property. To start with, recall the partial order of regions defined in Definition 5. Definition 15 defines Apriori property of slice transformation results w.r.t. this partial order:
Definition 15 (Apriori property of slice transformation results).
Let $\sigma$ be a metric computed by a slice transformation. The metric is said to satisfy the apriori property if for any two regions $\gamma_{1}\prec\gamma_{2}$, $\sigma[\gamma_{1}]\prec\sigma[\gamma_{2}]$ [^2] We use $\prec$ to denote a partial order for metrics, which may not be the linear order $<$ on real numbers., where $\sigma[\gamma]$ denotes the value of metric $\sigma$ on region $\gamma$.
Apriori property and early stopping. If a slice transformation computes a metric that satisfies the Apriori property, then filtering on the metric allows us to early stop. This can be best illustrated on the CRAWL for frequent itemset mining (both Section 3.1 and pipe 2). For frequent itemset mining, the metric ‘support’ satisfies the Apriori property. For example, for the standard frequent itemset mining described in Section 3.1, if we filter by ‘support > .1’, then if the region (Beer=1,) already has support <= .1, then we can stop there without considering any region (pattern) that is more fine-grained than (Beer=1,).
Degree-first evaluation and Apriori Algorithm for Frequent Itemset Mining. One can choose to leverage the Apriori property for early stopping in various ways. For example, one can do a degree-first333 the number of attributes in a region space is called the degree of the region space. evaluation of region spaces: We first consider evaluating all region spaces of degree 1 ([Beer], [Diaper], etc.), and then with the results of degree-1 regions, we filter by ‘support’, and only on those regions we consider region spaces of degree 2 ([Beer, Diaper], [Beer, Rum], etc.). This corresponds to a Breadth-First Search of regions. In fact, astute readers will recognize that this degree-first evaluation recovers the well known Apriori algorithm for Frequent Itemset Mining [Apr]. This shows that our systematic optimization from considering Apriori property captures such an algorithm as a very special case. And the same optimization applies to many other tasks, as long as we filter by a metric with the Apriori property and do degree-first evaluation w.r.t. region spaces.
Beyond degree-first evaluation. Clearly, besides degree-first evaluation, we also have other evaluation strategies, such as depth-first. An interesting strategy is optimistic region evaluation, where we evaluate all region spaces in parallel, but before doing the slice transformations on a region, we check whether the region can be skipped using the Apriori property. Selecting the optimal strategy for performance is crucial in implementing multi-relational algebra, but a detailed exploration is beyond this paper’s scope and will be addressed in future work.
4.3 Main application categories
@@
The multi-relational pipe service provides useful abstractions that benefit all these applications:
Simplified systems. In experimentation analysis platforms, teams often build custom slicing-analysis kernels, each involving tens of thousands of lines of code, typically using Apache Beam [Pyt] due to its flexibility over SQL. Our multi-relational algebra pipes streamline these into concise multi-relational pipes. These pipelines, sharing a relational data foundation with SQL, improve interoperability. The outcome is substantial simplification for both experimentation analysis and integration with relational databases.
Enhanced Capabilities. The ability to compose multi-relational operators has opened doors to powerful new functionalities. For instance, techniques such as “recurrent crawl” (Section 3.5) have leveraged these operators in system monitoring to achieve significantly better anomaly detection compared to traditional statistical methods alone. This showcases the power of multi-relational operators in expressing complex analysis patterns. Additionally, the flexibility of slice transformations allows for seamless integration with the Python Data Science stack, facilitating advanced analyses. In practice, this has enabled rapid identification of revenue anomalies and their correlation with campaign budgets, leading to substantial revenue increases in some cases.
Reduced Costs. Migrating ad-hoc experimentation analysis systems to multi-relational pipes offers a twofold advantage: it empowers these systems to analyze data at a more granular level and, counter-intuitively, leads to improved latency, scalability, and reduced resource costs over the ad-hoc constructions. This efficiency gain arises from the unified logical design of multi-relational algebra, which allows for systematic optimizations across a wide range of analytical tasks.
@@ Appendix A Emulating WINDOW functions
Appendix B Encoding a sparse relation
To avoid the waste of encoding 0 in Section 3.1, we observe that schemas are the only important thing in a multi-relation to tell whether we have multiple relations, or just a single relation. To explain this, recall that a transaction set can certainly be encoded succinctly as a set of tuples with different schemas, namely like
(Diaper=1,),
(Diaper=1, Beer=1,),
(Rum=1, Whiskey=1,),
Now, instead of directly encoding this as a multi-relation with different schemas (thus this multi-relation indeed contains multiple relations of different schemas), we can encode this as a single relation, by just declaring a single schema that is the list of all items. Note that, the tuples can still have different schemas, but encoding that there is a single schema makes this into a relation with one schema, instead of a multi-relation. We can visualize this relation as follows, note that there is no 0 any more, but values are just “missing” (we don’t even need to pad NULL).
Beer Diaper Rum Whisky Count
$\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{1}$
$\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{1}$
$\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{1}$
Table 3: Encoding a sparse table. We only use a wide schema to encode that this is a single relation.
Now, when we perform MultiSelect, the schema matching process will only find one relation as matching, since there is only one schema! Further, for a particular region space, the semantics is simply tuples that have values on that region space. For example, for region space [Diaper], the slice representation will simply identifies tuples that have value $1$ on Diaper (since $1$ is the only value in this relation), and therefore there is only a single region (Device=1), which always corresponds to a legit pattern. Clearly, ‘support‘ can be easily computed from here.
Appendix C Extended operators
@@
C.1 MultiGroupBy
In the RA, GroupBy partitions a relation into groups, and then uses an aggregation function to condense each group into a tuple, and finally we coalesce these tuples back into one relation. For Multi-Relational Algebra, the situation is similar, we partition a multi-relation into groups, where each group is a multi-relation. Then we transform each group into another multi-relation. Finally, we coalesce these multi-relations into one by flattening the group key into the regions of each multi-relation, and take union of all multi-relations.
@@ C.1 MultiGroupBy
MultiGroupBy is an operator as follows:
MultiGroupBy(
$\Psi$,
@@
    We partition $\Psi$ using by_attributes.
@@
    For each group, we apply multi_relational_transform to transform each multi-relation to another.
@@
    We flatten the group keys into the corresponding multi-relations, and take a union of all multi-relations.
C.2 CoMultiGroupBy
Consider two multi-relations $\Psi_{1}$ and $\Psi_{2}$, and we groupby for each on Country. Now, we want the multi-relation in $\Psi_{1}$ under Country=US to co-occur with the multi-relation in $\Psi_{2}$ under Country=US. Such a need of co-occurrence is not new: this also appears in Apache Beam CoGroupByKey [CoG]. However, unlike Beam which only needs to work with a data model of general dictionaries, here we must eventually output one multi-relation. Suppose we use $\Psi_{1}${Country=US} to denote the group Country=US after groupby, then co-groupby will put this together with $\Psi_{2}${Country=US}, so we have a pair of multi-relations ($\Psi_{1}${Country=US}, $\Psi_{2}${Country=US}). Then, we can use a MultiJoin to condense this pair of multi-relations back into one. After this, we are then in the situation of MultiGroupBy, which we know how to handle.
@@ C.2 CoMultiGroupBy
CoMultiGroupBy is an operator as follows:
CoMultiGroupBy(
$\Psi_{1}$,
$\Psi_{2}$,
@@
    We partition $\Psi_{1}$ and $\Psi_{2}$ by by_attributes to form groups, where each group is a multi-relation.
@@
    For each group key $K$, we have a multi-relation $\Psi_{1}\{K\}$, and another multi-relation $\Psi_{2}\{K\}$.
@@
    We join these two multi-relations into one using multi_join($\Psi_{1}\{K\}$, $\Psi_{2}\{K\}$).
After the above steps, we are in the same situation as MultiGroupBy, which we handle accordingly.